need to show workinga) Coffee Hut purchased a coffee machine on the 1 January 2017 for $58,000. The expected useful life is 7 years, or 100,000 c
Answer #1

1. Straight line method

Depreciation for 2017 = (58000-2000)/7 = $8000

Depreciation for 2018 = $8000

2. Reducing balance method

Depreciation rate = 7 years = 14.29%

Depreciation for 2017 = 58000*14.29% = $8288

Depreciation for 2018 = (58000-8288)*14.29% = $7104

3. Depreciation for 2017 = (58000-2000)*5000/100000 = $2800

Depreciation for 2018 = (58000-2000)*6500/100000 = $3640